The GOP deflator in Econoland is 200 on January 1, 2005. The deflator rises to 242 by January 1, 2007, and to 266.2 by January 1, 2008. n. What is the annual rate of inflation over the two-year period between January 1, 2005, and January 1, 2007?In other words, what constant yearly rate of infla- tion would lead to the price rise observed over those two years? b. What is the annual rate of inflation over the three- year period from January 1, 2005, to January 1, 2008? c. In general, if Po is the price level at the beginning of an n-year period, and P" is the price level at the end of that period, show that the annual rate of inflation It over that period satisfies the equation (1 + It)'' = P,/Po.


1
Expert's answer
2021-10-08T08:12:27-0400

a. constant yearly rate of inflation

"(1+pi)^2

=\frac{242}{200}"

"(\frac{242}{200})^\frac{1}{2} -1 = 10" %, "pi=10%" %

b. annual rate of inflation over the three-year period from January 1, 2005, to January 1, 2008

"(1+pi)^3=\frac{266.2}{200}"

"(\frac{266.2}{200})^\frac{1}{3} -1 = 10%" %

c. after 1 year price = "Po\times(1+pi)^n = Pn"

"(1+pi)^n = \frac{Pn}{Po}"